Science Report – 26 April 2018

1. Emergency EVAC EVA: No locations were scouted during either EVA today. Planning for an EVA dedicated to emergency procedures is in the works.

2. Sunspot and Solar Flare Monitoring: Our CGA continued his task of monitoring the sun for sunspot activity. The sunspots are getting smaller – they were not visible with the eye but could still be seen with the Skyris camera.

3. Shortwave Texting on EVA’s: Beartooth devices still not working. Hand signal communication was used instead.

4. GPS Route Measurements: GPS was taken on EVA 8 this morning which proved very helpful in re-locating the rover after exiting the moon valley area. A picture was taken of the GPS device when EVA 8 parked for exploration and the crew the followed the GPS back to that set of coordinates.

5. Waterless dish cleaning: No samples were taken today. No growth seen on samples from previous day. Previous day samples were put into the incubator in the science hab today to speed bacterial growth.

6. The MAG (Maximum Absorbency Garment) utilization study: No progress today. Possible that the continuation of this study from Crew 192 will be scrapped.

7. Spacesuit Helmet Fogging: Neither crew experienced fogging of the masks today. Both crews used a defogging spray and non-scratch cloth on the masks prior to their EVAs today. One crew member thought their mask had fogged, but it turned out to just be dirt.

8. Crew Comfort, Health and Safety on EVA: Measurements of the crew’s weight, blood pressure, SpO2 and pulse were taken before and after the EVA. Maximum heart rate was tracked by a mobile monitor during the EVA when the astronaut felt that they were working very hard (normally climbing a large hill). Gold 01 is able to track his heart rate over the entire course of the EVA thanks to mobile technology. Results can be exported and used in this study. Analysis will be done as the mission continues.

Copyright © The Mars Society. All rights reserved. | Main Site